OpenStudy (anonymous):

Determine the slope of the line that contains the points A(0,6) & B(4,0). Choose one answer. a. -5/4 b. undefined c. 4/5 d. -3/2

OpenStudy (aroub):

slope formula: y2-y1/x2-x1

OpenStudy (aroub):

=> 0-6/4-0 = -6/4=-3/2

OpenStudy (anonymous):

d

OpenStudy (anonymous):

what she said ^^